# Scaling Up to Excellence: Practicing Model Scaling for Photo-Realistic Image Restoration In the Wild

### v1

@@ 1 Introduction
With the development of image restoration (IR), expectations for the perceptual effects and intelligence of IR results have significantly increased. IR methods based on generative priors [42, 82, 67, 49] leverage powerful pre-trained generative models to introduce high-quality generation and prior knowledge into IR, bringing significant progress in these aspects. Continuously enhancing the capabilities of the generative prior is key to achieving more intelligent IR results, with model scaling being a crucial and effective approach. There are many tasks that have obtained astonishing improvements from scaling, such as SAM [44] and large language models [7, 74, 73]. This further motivates our effort to build large-scale, intelligent IR models capable of producing ultra-high-quality images. However, due to engineering constraints such as computing resources, model architecture, training data, and the cooperation of generative models and IR, scaling up IR models is challenging.
@@
In this work, we introduce SUPIR (Scaling-UP IR), the largest-ever IR method, aimed at exploring greater potential in visual effects and intelligence. Specifically, SUPIR employs StableDiffusion-XL (SDXL) [63] as a powerful generative prior, which contains 2.6 billion parameters. To effectively apply this model, we design and train a adaptor with more than 600 million parameters. Moreover, we have collected over 20 million high-quality, high-resolution images to fully realize the potential offered by model scaling. Each image is accompanied by detailed descriptive text, enabling the control of restoration through textual prompts. We also utilize a 13-billion-parameter multi-modal language model to provide image content prompts, greatly improving the accuracy and intelligence of our method. The proposed SUPIR model demonstrates exceptional performance in a variety of IR tasks, achieving the best visual quality, especially in complex and challenging real-world scenarios. Additionally, the model offers flexible control over the restoration process through textual prompts, vastly broadening the possibility of IR. Fig. 1 illustrates the effects by our model, showcasing its superior performance.
@@
Our work goes far beyond simply scaling. While pursuing an increase in model scale, we face a series of complex challenges. First, when applying SDXL for IR, existing Adaptor designs either too simple to meet the complex requirements of IR [59] or are too large to train together with SDXL [95]. To solve this problem, we trim the ControlNet and designed a new connector called ZeroSFT to work with the pre-trained SDXL, aiming to efficiently implement the IR task while reducing computing costs. In order to enhance the model’s ability to accurately interpret the content of low-quality images, we fine-tune the image encoder to improve its robustness to variations in image degradation. These measures make scaling the model feasible and effective, and greatly improve its stability. Second, we amass a collection of 20 million high-quality, high-resolution images with descriptive text annotations, providing a solid foundation for the model’s training. We adopt a counter-intuitive strategy by incorporating poor quality, negative samples into training. In this way, we can use negative quality prompts to further improve visual effects. Our results show that this strategy significantly improves image quality compared to using only high-quality positive samples. Finally, powerful generative prior is a double-edged sword. Uncontrolled generation may reduce restoration fidelity, making IR no longer faithful to the input image. To mitigate this low-fidelity issue, we propose a novel restoration-guided sampling method. All these strategies, coupled with efficient engineering implementation, are key to enabling the scaling up of SUPIR, pushing the boundaries of advanced IR. This comprehensive approach, encompassing everything from model architecture to data collection, positions SUPIR at the forefront of image restoration technology, setting a new benchmark for future advancements.
@@ 2 Related Work
The goal of IR is to convert degraded images into high-quality degradation-free versions [91, 26, 98, 99, 89, 22]. In the early stage, researchers independently explored different types of image degradation, such as super-resolution (SR) [19, 20, 13], denoising [90, 92, 11], and deblurring [14, 60, 72]. However, these methods are often based on specific degradation assumptions [50, 58, 25] and therefore lack generalization ability to other degradations [29, 97, 53]. Over time, the need for blind restoration methods that are not based on specific degradation assumptions has grown [78, 34, 47, 5, 35, 48, 46, 94, 10]. In this trend, some methods [93, 81] approximate synthesize real-world degradation by more complex degradation models, and are well-known for handling multiple degradation with a single model. Recent research, such as DiffBIR [49], unifies different restoration problems into a single model. In this paper, we adopt a similar setting to DiffBIR and use a single model to achieve effective processing of various severe degradations.
@@
Generative priors are adept at capturing the inherent structures of the image, enabling the generation of images that follow natural image distribution. The emergence of GANs [23, 64, 40, 39] has underscored the significance of generative priors in IR. Various approaches employ these priors, including GAN inversion [27, 57, 2, 62, 4], GAN encoders [103, 9], or using GAN as the core module for IR [87, 80]. Beyond GANs, other generative models can also serve as priors [75, 102, 101, 10, 36, 55, 100]. Our work primarily focuses on generative priors derived from diffusion models [31, 70, 61, 67, 65, 71], which excel in controllable generation [59, 95, 15, 18, 32] and model scaling [63, 66, 68]. Diffusion models have also been effectively used as generative priors in IR [42, 82, 49, 67, 77]. However, these diffusion-based IR methods’ performance is constrained by the scale of the used generative models, posing challenges in further enhancing their effectiveness.
@@
is an important means to further improve the capabilities of deep-learning models. The most typical examples include the scaling of language models [7, 74, 73], text-to-image generation models [68, 67, 63, 12, 85, 37], and image segmentation models [44]. The scale and complexity of these models have increased dramatically, with billions or even hundreds of billions of parameters, but these parameters also lead to extraordinary performance improvements, demonstrating the potential of model scaling [38]. However, scaling up is a systematic problem, involving model design, data collection, computing resources, and other limitations. Many other tasks have not yet been able to enjoy the substantial performance improvements brought by scaling up. IR is one of them.
@@ 3 Method
An overview of the proposed SUPIR method is shown in Fig. 2. We introduce our method from three aspects: Sec. 3.1 introduces our network designs and training method; Sec. 3.2 introduces the collection of training data and the introduction of text modality; and Sec. 3.3 introduces the diffusion sampling method for image restoration.
@@ 3.1 Model Scaling Up
There are not many choices for the large-scale generative models. The only ones to consider are Imagen [68], IF [16], and SDXL [63]. Our selection settled on SDXL for the following reasons. Imagen and IF prioritize text-to-image generation and rely on a hierarchical approach. They first generate small-resolution images and then hierarchically upsample them. SDXL directly generates a high-resolution image without hierarchical design, which is more aligned with our objectives, as it utilizes its parameters effectively for image quality improvement rather than text interpretation. Additionally, SDXL employs a Base-Refine strategy. In the Base model, diverse but lower-quality images are generated. Subsequently, the Refine model enhances the perceptual quality of these images. Compared to the Base model, the Refine model uses training images with significantly higher quality but less diverse. Considering our strategy to train with an extensive dataset of high-quality images, the two-phase design of SDXL becomes superfluous for our needs. We opt for the Base model, which has a greater number of parameters, making it an ideal backbone for our generative prior.
@@
Considering the SDXL model as our chosen prior, we need an adaptor that can steer it to restore images according to the provided LQ inputs. The Adaptor is required to identify the content in the LQ image and to finely control the generation at the pixel level. LoRA [32], T2I adaptor [59], and ControlNet [95] are existing diffusion model adaptation methods, but none of them meet our requirements: LoRA limits generation but struggles with LQ image control; T2I lacks capacity for effective LQ image content identification; and ControlNet’s direct copy is challenging for the SDXL model scale. To address this issue, we design a new adaptor with two key features, as shown in Fig. 3(a). First, we keep the high-level design of ControlNet but employ network trimming [33] to directly trim some blocks within the trainable copy, achieving an engineering-feasible implementation. Each block within the encoder module of SDXL is mainly composed of several Vision Transformer (ViT) [21] blocks. We identified two key factors contributing to the effectiveness of ControlNet: large network capacity and efficient initialization of the trainable copy. Notably, even partial trimming of blocks in the trainable copy retains these crucial characteristics in the adaptor. Therefore, we simply trim half of the ViT blocks from each encoder block, as shown in Fig. 3(b). Second, we redesign the connector that links the adaptor to SDXL. While SDXL’s generative capacity delivers excellent visual effects, it also renders pixel-level precise control challenging. ControlNet employs zero convolution for generation guidance, but relying solely on residuals is insufficient for the control required by IR. To amplify the influence of LQ guidance, we introduced a ZeroSFT module, as depicted in Fig. 3(c). Building based on zero convolution, ZeroSFT encompasses an additional spatial feature transfer (SFT) [79] operation and group normalization [84].
@@ 3.2 Scaling Up Training Data
The scaling of the model requires a corresponding scaling of the training data [38]. But there is no large-scale high-quality image dataset available for IR yet. Although DIV2K [3] and LSDIR [1] offer high image quality, they are limited in quantity. Larger datasets like ImageNet (IN) [17], LAION-5B[69], and SA-1B [44] contain more images, but their image quality does not meet our high standards. To this end, we collect a new large-scale dataset of high-resolution images, which includes 20 million 1024$\times$1024 high-quality, texture-rich, and content-clear images. A comparison on scales of the collected dataset and the existing dataset is shown in Fig. 3. We also included an additional 70K unaligned high-resolution facial images from FFHQ-raw dataset [40] to improve the model’s face restoration performance. In Fig. 5(a), we show the relative size of our data compared to other well-known datasets.
@@
Diffusion models are renowned for their ability to generate images based on textual prompts. We believe that textual prompts can also significantly aid IR for the following reasons: (1) Understanding image content is crucial for IR. Existing frameworks often overlook or implicitly handle this understanding [24, 29]. By incorporating textual prompts, we explicitly convey the understanding of LQ images to the IR model, facilitating targeted restoration of missing information. (2) In cases of severe degradation, even the best IR models may struggle to recover completely lost information. In such cases, textual prompts can serve as a control mechanism, enabling targeted completion of missing information based on user preferences. (3) We can also describe the desired image quality through text, further enhancing the perceptual quality of the output. See Fig. 1(b) for some examples. To this end, we make two main modifications. First, we revise the overall framework to incorporate the LLaVA multi-modal large language model [52] into our pipeline, as shown in Fig. 2. LLaVA takes the degradation-robust processed LQ images $x_{LQ}^{\prime}=\mathcal{D}(\mathcal{E}_{\mathrm{dr}}(x_{LQ}))$ as input and explicitly understands the content within the images, outputting in the form of textual descriptions. These descriptions are then used as prompts to guide the restoration. This process can be automated during testing, eliminating the need for manual intervention. Secondly, following the approach of PixART [12], we also collect textual annotations for all the training images, to reinforce the role of textual control during the training of out model. These two changes endow SUPIR with the ability to understand image content and to restore images based on textual prompts.
@@ 4.1 Model Training and Sampling Settings
For training, the overall training data includes 20 million high-quality images with text descriptions, 70K face images and 100K negative-quality samples and the corresponding negative prompts. To enable a larger batch size, we crop them into 512$\times$512 patches during training. We train our model using a synthetic degradation model, following the setting used by Real-ESRGAN [81], the only difference is that we resize the produced LQ images to 512$\times$512 for training. We use the AdamW optimizer [54] with a learning rate of 0.000010.000010.00001. The training process spans 10 days and is conducted on 64 Nvidia A6000 GPUs, with a batch size of 256. For testing, the hyper-parameters are $T$=100, $\lambda_{\mathrm{cfg}}$=7.5, and $\tau_{r}=4$. Our method is able to process images with the size of 1024$\times$1024. We resize the short side of the input image to 1024 and crop a 1024$\times$1024 sub-image for testing, and then resize it back to the original size after restoration. Unless stated otherwise, prompts will not be provided manually – the processing will be entirely automatic.
@@ 4.2 Comparison with Existing Methods
Our method can handle a wide range of degradations, and we compare it with the state-of-the-art methods with the same capabilities, including BSRGAN [93], Real-ESRGAN [81], StableSR [77], DiffBIR [49] and PASD [88]. Some of them are constrained to generating images of 512$\times$512 size. In our comparison, we crop the test image to meet this requirement and downsample our results to facilitate fair comparisons. We conduct comparisons on both synthetic data and real-world data.
@@
Single: SR (×4$\times$ 4× 4) BSRGAN 25.06 0.6741 0.2159 0.2214 0.6169 70.38
@@
Single: SR (×8$\times$ 8× 8) BSRGAN 22.26 0.5212 0.3523 0.2069 0.5836 67.04
@@
Mixture: Blur ($\sigma$=2) $+$ SR (×4$\times$ 4× 4) BSRGAN 24.97 0.6572 0.2261 0.2127 0.5984 69.44
@@
Mixture: SR (×4$\times$ 4× 4)$+$ Noise ($\sigma$=40) BSRGAN 17.74 0.3816 0.5659 0.1006 0.4166 51.25
@@
Mixture: Blur ($\sigma$=2) $+$ SR (×4$\times$ 4× 4)$+$ Noise ($\sigma$=20)$+$ JPEG ($q$=50) BSRGAN 22.88 0.5397 0.3445 0.1838 0.5402 64.81

### v2

@@ 1 Introduction
The development of image restoration (IR) has greatly elevated expectations for both the perceptual effects and the intelligence of IR results. IR methods based on generative priors [42, 82, 67, 49] leverage powerful pre-trained generative models to introduce high-quality generation and prior knowledge into IR, bringing significant progress in these aspects. Continuously improving the capabilities of the generative prior is key to achieving better IR results, with model scaling being a crucial and effective approach. There are many tasks that have obtained astonishing improvements from scaling, such as SAM [44] and large language models (LLMs) [7, 74, 73]. This further drives our pursuit of constructing large-scale, intelligent IR models that can produce ultra-high-quality images. However, due to engineering constraints such as computing resources, model architecture, training data, and the cooperation of generative models and IR, scaling up IR models is challenging.
@@
In this work, we introduce SUPIR (Scaling-UP IR), the largest-ever IR method, aimed at exploring greater potential in restoration visual effects and intelligence. Specifically, SUPIR employs StableDiffusion-XL (SDXL) [63] as a powerful generative prior, which contains 2.6 billion parameters. To effectively deploy this model in IR, we design and train a large-scale adaptor that incorporates a novel component named the ZeroSFT connector. To maximize the benefits of model scaling, we collect a dataset of over 20 million high-quality, high-resolution images, each accompanied by detailed descriptive text. We utilize a 13-billion-parameter multi-modal language model to provide image content prompts, greatly improving the accuracy and intelligence of our method. The proposed SUPIR model demonstrates exceptional performance in a variety of IR tasks, achieving the best visual quality, especially in complex and challenging real-world scenarios. Additionally, the model offers flexible control over the restoration process through textual prompts, vastly broadening the possibility of IR. Fig. 1 illustrates the effects by our model.
@@
Our work goes far beyond simply scaling. While pursuing an increase in model scale, we face a series of complex challenges. First, existing adaptor designs either too simple to meet the complex requirements of IR [59] or are too large to train together with SDXL [95]. To solve this problem, we trim the ControlNet and designed a new connector called ZeroSFT to work with the pre-trained SDXL, aiming to efficiently implement the IR task while reducing computing costs. In order to enhance the model’s ability to accurately interpret the content of low-quality images, we fine-tune the image encoder to improve its robustness to variations in image degradation. These measures make scaling the model feasible and effective, and greatly improve its stability. Second, we collect 20 million high-quality, high-resolution images with descriptive text annotations, providing a solid foundation for the model’s training. We employ a counter-intuitive approach by integrating poor-quality samples into our training process. This allows us to enhance visual effects by utilizing prompts to guide the model away from negative qualities. Finally, powerful generative prior is a double-edged sword. Uncontrolled generation may reduce restoration fidelity, making IR no longer faithful to the input image. To address the issue of low fidelity, we introduce the concept of restoration-guided sampling. By integrating these strategies with efficient engineering practices, we not only facilitate the scaling up of SUPIR but also push the frontiers of advanced IR.
@@ 2 Related Work
The goal of IR is to convert degraded images into high-quality degradation-free images [91, 26, 98, 99, 89, 22]. In the early stage, researchers independently explored different types of image degradation, such as super-resolution (SR) [19, 20, 13], denoising [90, 92, 11], and deblurring [14, 60, 72]. However, these methods are often based on specific degradation assumptions [50, 58, 25] and therefore lack generalization ability to other degradations [29, 97, 53, gu2024interpretability]. Over time, the need for blind IR methods that are not based on specific degradation assumptions has grown [25, 34, 47, 5, 35, 48, 46, 94, 10]. In this trend, some methods [93, 81] synthesize real-world degradation by more complex degradation models, and are well-known for handling multiple degradation with a single model. DiffBIR [49] unifies different restoration problems into a single model. In this paper, we adopt a similar setting to DiffBIR and use a single model to achieve effective processing of various severe degradations.
@@
Generative priors are adept at capturing the inherent structures of the image, enabling the generation of images that follow natural image distribution. The emergence of GANs [23, 64, 40, 39] has underscored the significance of generative priors in IR. Various approaches employ generative priors, including GAN inversion [27, 57, 2, 62, 4], GAN encoders [103, 9], or using GAN as the core module for IR [87, 80]. Beyond GANs, other generative models can also serve as priors [75, 102, 101, 10, 36, 55, 100]. Our work primarily focuses on generative priors derived from diffusion models [31, 70, 61, 67, 65, 71], which excel in controllable generation [59, 95, 15, 18, 32] and model scaling [63, 66, 68]. Diffusion models have also been effectively used as generative priors in IR [42, 82, 49, 67, 77]. However, these diffusion-based IR methods’ performance is constrained by the scale of the used generative models, posing challenges in further enhancing their effectiveness.
@@
is an important means to further improve the capabilities of deep-learning models. The most typical examples include the scaling of language models [7, 74, 73], text-to-image generation models [68, 67, 63, 12, 85, 37], and image segmentation models [44]. The scale and complexity of these models have increased dramatically, now encompassing billions of parameters. This increase in parameters has also resulted in significant performance enhancements, showcasing the immense potential of model scaling [38]. However, scaling up is a systematic problem, involving model design, data collection, computing resources, and other limitations. Many other tasks have not yet been able to enjoy the substantial performance improvements brought by scaling up. IR is one of them.
@@ 3 Method
An overview of the proposed SUPIR method is shown in Fig. 2. We introduce our method from three aspects: Sec. 3.1 introduces our network designs and training method; Sec. 3.2 introduces the collection of training data and the introduction of textual modality; and Sec. 3.3 introduces the diffusion sampling method for IR.
@@ 3.1 Model Scaling Up
There are not many choices for the large-scale generative models. The only ones to consider are Imagen [68], IF [16], and SDXL [63]. Our selection settled on SDXL for the following reasons. Imagen and IF prioritize text-to-image generation and rely on a hierarchical approach. They first generate small-resolution images and then hierarchically upsample them. SDXL aligns with our objectives by directly generating high-resolution images without a hierarchical design, effectively using its parameters to improve image quality rather than focusing on text interpretation. Additionally, SDXL employs a Base-Refine strategy. In the Base model, diverse but lower-quality images are generated. Subsequently, the Refine model, utilizing training images of significantly higher quality but lesser diversity than those used by the Base model, enhances the images’ quality. Given our approach of training with a vast dataset of high-quality images, the dual-phase design of SDXL becomes redundant for our objectives. We opt for the Base model, which has a greater number of parameters, making it an ideal generative prior.
@@
Considering the SDXL model as our chosen prior, we need an adaptor that can steer it to restore images according to the provided LQ inputs. The Adaptor is required to identify the content in the LQ image and to finely control the generation at the pixel level. LoRA [32], T2I adaptor [59], and ControlNet [95] are existing diffusion model adaptation methods, but none of them meet our requirements: LoRA limits generation but struggles with LQ image control; T2I lacks capacity for LQ image content identification; and ControlNet’s direct copy is challenging for the SDXL model scale. To address this issue, we design a new adaptor with two key features, as shown in Fig. 3(a). First, we keep the high-level design of ControlNet but employ network trimming [33] to directly trim some blocks within the trainable copy, achieving an engineering-feasible implementation. Each block within the encoder module of SDXL is mainly composed of several Vision Transformer (ViT) [21] blocks. We identified two key factors contributing to the effectiveness of ControlNet: large network capacity and efficient initialization of the trainable copy. Notably, even partial trimming of blocks in the trainable copy retains these crucial characteristics in the adaptor. Therefore, we simply trim half of the ViT blocks from each encoder block, as shown in Fig. 3(b). Second, we redesign the connector that links the adaptor to SDXL. While SDXL’s generative capacity delivers excellent visual effects, it also renders pixel-level control challenging. ControlNet employs zero convolution for generation guidance, but relying solely on residuals is insufficient for the control required by IR. To amplify the influence of LQ guidance, we introduced a ZeroSFT module, as depicted in Fig. 3(c). Building based on zero convolution, ZeroSFT encompasses an additional spatial feature transfer (SFT) [79] operation and group normalization [84].
@@ 3.2 Scaling Up Training Data
The scaling of the model requires a corresponding scaling of the training data [38]. But there is no large-scale high-quality image dataset available for IR yet. Although DIV2K [3] and LSDIR [1] offer high image quality, they are limited in quantity. Larger datasets like ImageNet (IN) [17], LAION-5B [69], and SA-1B [44] contain more images, but their image quality does not meet our high standards. To this end, we collect a large-scale dataset of high-resolution images, which includes 20 million 1024$\times$1024 high-quality, texture-rich images. A comparison on the scales of the collected dataset and the existing dataset is shown in Fig. 3. We also included an additional 70K unaligned high-resolution facial images from the FFHQ-raw dataset [40] to improve the model’s face restoration performance. In Fig. 5(a), we show the relative size of our data compared to other well-known datasets.
@@
Diffusion models are renowned for their ability to generate images based on textual prompts. We believe that textual prompts can also aid IR: (1) Understanding image content is crucial for IR. Existing frameworks often overlook or implicitly handle this understanding [24, 29]. By incorporating textual prompts, we explicitly convey the understanding of LQ images to the IR model, facilitating targeted restoration of missing information. (2) In cases of severe degradation, even the best IR models struggle to recover completely lost information. In such cases, textual prompts can serve as a control mechanism, enabling targeted completion of missing information based on user preferences. (3) We can also describe the desired image quality through text, further enhancing the perceptual quality of the output. See Fig. 1(b) for some examples. To this end, we make two main modifications. First, we revise the overall framework to incorporate the LLaVA multi-modal LLM [52] into our pipeline, as shown in Fig. 2. LLaVA takes the degradation-robust processed LQ images $x_{LQ}^{\prime}=\mathcal{D}(\mathcal{E}_{\mathrm{dr}}(x_{LQ}))$ as input and explicitly understands the content within the images, outputting in the form of textual descriptions. These descriptions are then used as prompts to guide the restoration. This process can be automated during testing, eliminating the need for manual intervention. Secondly, following the approach of PixART [12], we also collect textual annotations for all the training images, to reinforce the role of textual control during the training of out model. These two changes endow SUPIR with the ability to understand image content and to restore images based on textual prompts.
@@ 4.1 Model Training and Sampling Settings
For training, the overall training data includes 20 million high-quality images with text descriptions, 70K face images and 100K negative-quality samples, together their corresponding prompts. To enable a larger batch size, we crop images into 512$\times$512 patches during training. We train our model using a synthetic degradation model, following the setting used by Real-ESRGAN [81], the only difference is that we resize the produced LQ images to 512$\times$512 for training. We use the AdamW optimizer [54] with a learning rate of $0.00001$. The training process spans 10 days and is conducted on 64 Nvidia A6000 GPUs, with a batch size of 256. For testing, the hyper-parameters are $T$=100, $\lambda_{\mathrm{cfg}}$=7.5, and $\tau_{r}=4$. Our method is able to process images with the size of 1024$\times$1024. We resize the short side of the input image to 1024 and crop a 1024$\times$1024 sub-image for testing, and then resize it back to the original size after restoration. Unless stated otherwise, prompts will not be provided manually – the processing will be entirely automatic.
@@ 4.2 Comparison with Existing Methods
Our method can handle a wide range of degradations, and we compare it with the latest methods with the same capabilities, including BSRGAN [93], Real-ESRGAN [81], StableSR [77], DiffBIR [49] and PASD [88]. Some of them are constrained to generating images of 512$\times$512 size. In our comparison, we crop the test image to meet this requirement and downsample our results. We conduct comparisons on both synthetic data and real-world data.
@@
Single: SR (×$\times 4$) BSRGAN 25.06 0.6741 0.2159 0.2214 0.6169 70.38
@@
Single: SR (×$\times 8$) BSRGAN 22.26 0.5212 0.3523 0.2069 0.5836 67.04
@@
Mixture: Blur ($\sigma$=2) $+$ SR (×$\times 4$) BSRGAN 24.97 0.6572 0.2261 0.2127 0.5984 69.44
@@
Mixture: SR (×$\times 4$)$+$ Noise ($\sigma$=40) BSRGAN 17.74 0.3816 0.5659 0.1006 0.4166 51.25
@@
Mixture: Blur ($\sigma$=2) $+$ SR (×$\times 4$)$+$ Noise ($\sigma$=20)$+$ JPEG ($q$=50) BSRGAN 22.88 0.5397 0.3445 0.1838 0.5402 64.81